# Matrix product state approach to lossy boson sampling and noisy IQP sampling

### v1

@@ IV.3 Entanglement Rényi entropy of the output state
Using those tools, we now evaluate the ERE between the bipartite subsystems of the same size, where the ERE is maximum [48, 49]; therefore, $l$ is chosen to be $\lfloor N/2\rfloor$. To do that, we now compute the eigenvalues of the reduced density matrix on the subsystem u (or equivalently the subsystem d).
@@
where $\kappa>0$ is a constant, the ERE is lower bounded by $\Omega(N^{\kappa})$. Hence, by choosing $\kappa\to 0$, i.e., $\eta$ scales faster than $1/\sqrt{N}$, the corresponding lossy boson sampling is hard to simulate using our MPS method. Therefore, the approximability achieved by our MPS method exhibits a transition at $\eta=\Theta(1/\sqrt{N})$, which is consistent with the GBS case [33].
@@
We now compare our results with those of previous studies [25, 24]. In the previous work, classical algorithms are proposed whose classically simulable range of $\eta$ is $\eta=O(1/\sqrt{N})$. The algorithms employed a method that approximates the lossy input state as a state that is easy to classically simulate, such as the closest thermal state or the closest separable state. An obvious crucial limitation of this approach is that the closest thermal state or the closest separable state to a given lossy input state remains fixed once the parameters of the lossy boson sampling are determined. That means that when the system is not sufficiently lossy, no thermal state or separable state can approximate the lossy input state with adequate accuracy. Furthermore, the simulation accuracy cannot be improved by allocating additional computational resources because the closest easy-to-simulate state is fixed.
@@
Compared to that, although our MPS algorithm exhibits a similar scaling in an asymptotic regime as $\eta<O(1/\sqrt{N})$, i.e., our MPS method might not provide an advantage over the previously known classical algorithm asymptotically, it shows a wider simulable range than the previous work in a non-asymptotic regime, which is more practically relevant, because even if the loss rate is not high enough, our algorithm always runs and can achieve a target accuracy by using more computational resources, i.e., by increasing the bond dimension. We emphasize that this improvement significantly enhances simulation performance in simulating lossy GBS [26, 33, 50]. To be more specific, Ref. [26] shows that the closest nonclassical state renders a similar classically simulable range of $\eta$, $\eta=O(1/\sqrt{N})$, which, however, was insufficient to simulate the experimental GBS. Nonetheless, Ref. [33] developed a classical algorithm based on MPS, just like ours, and could simulate the state-of-the-art experimental GBS, which clearly indicates that the MPS method significantly outperforms other methods in practice even though the asymptotic scaling is similar.
@@
We now compare this with another previous result based on MPO, which numerically finds a similar classically simulable range $\eta=O(1/\sqrt{N})$ [27]. Although the previous result yields the same strength in the sense that the accuracy can be improved by adding more computational resources, our work demonstrates a significant enhancement of the numerical simulability range. More specifically, the computational time cost of MPS and MPO simulations in a circuit of depth $D$ is $T=O(MDd^{4}\chi^{3})$ and $T=O(MDd^{8}\chi^{3})$, respectively. In terms of memory requirement, the memory cost of MPS is $O(\chi^{2}dM)$, while that of MPO is $O(\chi^{2}d^{2}M)$. In the MPO case, the local Hilbert space dimension $d$ is the square of that in the MPS case. This can lead to a higher computational cost when $d=N+1$ and $N$ is sufficiently large, as is the case in the boson sampling scenario. More crucially, as mentioned earlier, classical correlation can contribute to the increase in the bond dimension of MPO, whereas it cannot for MPS. Therefore, the bond dimensions between the two methods also differ between MPS and MPO. For example, the MPO method requires a bond dimension of the order of $10^{4}$, even for an input photon of less than $10$, with a loss rate of $0.7$. Our MPS method, on the other hand, requires only a bond dimension of the order of $10$ (note that the definitions of approximation errors differ, but a significant difference in bond dimension is evident).
@@
We also justify this commutation relation without the Haar-random assumption. It has been known that in the lossless case, the linear-optical circuit that renders the maximum entanglement entropy is the one that transforms the initial bosonic creation operators as [33, 35]
@@ IV.4 Numerical analysis of the computational cost
To understand how large parameter regimes our classical algorithm can simulate in practice, we now directly estimate the bond dimension for simulating the boson sampling algorithm with the photon number $N$ and transmission rate $\eta$, with a fixed MPS approximation error $\epsilon$ of $0.01$. Our MPS approximation error is defined by the sum of the discarded eigenvalues when the system is bipartitioned into two halves. Before we present our results, we emphasize that for certain regimes, other existing methods, such as the Clifford-Clifford algorithm, may outperform our method [23, 39]. Hence, one should interpret our results as a boundary set by our MPS method, rather than the boundary for the classically simulable regime for all possible methods.
In this numerical analysis, we consider three different families of circuits: (a) the worst-case construction, (b) the Haar-random unitary, and (c) an experimentally motivated structure. The purpose of this comparison is to examine how the required bond dimension varies across different choices of circuit unitaries. Here, the worst-case construction has $\theta_{j}=\pi/4$ for all $j$ to produce the largest amount of entanglement between the bipartitions we consider, and the experimentally motivated structure follows the same structure implemented in Ref. [12]. For (c), although the experiment in the reference implements GBS, we constructed a circuit unitary from this setup. Since the number of modes of this unitary is restricted as $M=2k^{2}$ for an integer $k$, for a given photon number $N$, we set $k$ as the nearest integer to $N/\sqrt{2}$ so that $M=2k^{2}$ is the closest value to $N^{2}$. However, we inject $\lceil N/2\rceil$ photons into the middle of the modes of the first half, and the remaining $\lfloor N/2\rfloor$ photons into the middle of the modes of the second half. This is because the whole unitary in this case is constructed by combining two unitaries of half the size; inputting the same number of single photons in both bipartitions and inputting single photons in the middle can help maximize the entanglement.
We present the numerical results in Fig. 6. First of all, Fig. 6(a) presents the required bond dimension in the worst-case circuit. Fig. 6(b) is when the unitary of the boson sampling circuit is drawn from a Haar random. The required resource in this result is comparable to the worst-case scenario, indicating that highly entangled configurations are typical in Haar random cases. One can see that if the loss rate is approximately $50\%$, which is the case for GBS experiments in Refs. [13, 15, 17], and the input photon number is $40$, the required bond dimension is $10^{7}$, which is already quite large to implement in practice.
@@
Figure 6: Required bond dimension and memory to achieve an MPS approximation error $\epsilon=0.01$. Panels (a)–(c) show the required bond dimension, and panels (d)–(f) show the required memory. (a) Required bond dimension where the unitary is in the worst-case scenario. This implies the highest cost for a specific input photon number and photon loss. (b) Required bond dimension when the Haar random unitary is a unitary corresponding to the circuit in boson sampling. (c) Required bond dimension when a circuit unitary is the same structure as the GBS experiment of USTC [12]. (d) Required memory where the unitary is in the worst-case scenario. (e) Required memory when the Haar random unitary is a unitary corresponding to the circuit in boson sampling. (f) Required memory when a circuit unitary is the same structure as the GBS experiment of USTC [12].
According to the required bond dimension, the memory complexity of the circuit can be estimated as shown in Fig. 6(d)–(f) [33]. Assuming that each complex number requires 8 bytes, the memory usage can be estimated as the total number of elements multiplied by 8 bytes:
@@ IV.6 Cat-state boson sampling
respectively. Note that as the amplitude of the cat state $\gamma\rightarrow 0$, the odd (even) cat state approaches $|1\rangle$ ($|0\rangle$), thereby recovering single-photon inputs. Moreover, cat state sampling is also known to be hard to classically simulate [51]. The circuit has the same structure as before, except that the input of the first $N$ mode is an odd cat state. The photon loss can be gathered in front of the circuit as in Fig. 1, and the cat state input goes through the total photon loss channel and transforms to the following state,
@@ V.2 Entanglement Rényi entropy of the output state
Hence, it shows that our MPS method is efficient up to the point where the depth scales logarithmically with the number of qubits $n$. Notably, this noise threshold is consistent with a recent result in Ref. [42], which employs a completely different method.
Figure 9: Simulation of depolarizing noise in IQP circuit. (a) Pauli $Z$ noises can commute with the very first gate of the circuit, as it is diagonal in the computational basis. (b) The noises are sampled into Pauli gates, where each $X$ and $Y$ gate occurs with probability $q$. And the input $|+\rangle$ goes through the Pauli $Z$ noise channel and becomes a mixed state $\hat{\tau}$. (c) We sample from an ensemble decomposition of the mixed state, drawing $|\phi_{+}\rangle$ or $|\phi_{-}\rangle$ with equal probability $1/2$. The $Y$ gates can be decomposed into $X$ and $Z$ gates. Here, because IQP gates are diagonal in the $Z$ basis, we absorb all $Z$-errors into the ideal IQP layer; in contrast, $X$-errors do not commute and therefore remain as the explicit noise layer.
@@ V.3 Generalization into depolarizing noise
Since the Rényi entropy of the state in a diagonal form depends only on the coefficients, the ERE of this state is exactly the same as that we obtained from the previous section. Therefore, as in the dephasing noise case, when the circuit depth $d\geq O(\log n/|\log(1-2q)|)=O(\log n/|\log(1-2p)|)$, the IQP sampling under depolarizing noise is efficiently classically simulable. When the noise rate is small, this depth approaches $O(\frac{1}{p}\log n)$, which corresponds to the critical depth in Ref. [42].
@@ V.4 Numerical analysis of noisy IQP sampling
Since we always assume the worst-case IQP circuit regardless of the circuit depth, even with $d=1$, a highly complicated IQP circuit is considered. Due to this reason, the required bond dimension is the maximum under the same number of qubits and the same noise rate per qubit and gate. Obviously, in practice, the complexity of an IQP circuit depends on its circuit depth [54, 45, 30, 55]. Implementing a highly complicated IQP circuit in a very shallow depth is demanding in practice; thus, we expect that the bond dimension for extremely low-depth cases overestimates the complexity. In other words, for an extremely low depth, though the required bond dimension is shown to be high in our figure, the corresponding IQP sampling can be classically efficiently simulated because entanglement cannot spread to the overall qubits. Except for that case, when the circuit depth increases, the overall noise rate also increases; therefore, the complexity of the simulation decreases.
For IQP sampling, the previous work shows the simulability beyond the critical depth $d_{c}>\Theta(p^{-1}\log(p^{-1}))$ [46]. According to the reference, the phase transition for percolation occurs at depths $d_{c}\approx 33$ for noise rate $p=0.05$ and $d_{c}\approx 117$ for $p=0.02$. However, since the IQP circuit is classically simulable using MPS with a bond dimension of order $10^{4}$, which may still be possible using a large classical computer, our MPS algorithm can cover a much wider range; more specifically, ours can simulate an IQP circuit with lower depth than the critical depth of the previous work. For the calculation of simulable memory of IQP, we use the same memory resource formula as in boson sampling.
@@ VI Discussion and conclusion
We remark that although our method for lossy boson sampling covers Fock-state and cat-state inputs, whether it can be further generalized to an arbitrary input state is still open. Also, while we assumed a conventional IQP circuit [19], a recent study in Ref. [42] extended the IQP circuit to be augmented by CNOT gates. Our method does not immediately extend to those IQP circuits yet. It is open to improve our algorithm to cover a broader class of quantum circuits. Also, we assumed a lossy channel or Pauli noise for boson sampling and IQP sampling, respectively. Generalizing our method to cover more general noise channels would be an important open question.

### v2

@@ IV.3 Entanglement Rényi entropy of the output state
Using those tools, we now evaluate the ERE between the bipartite subsystems of the same size, where the ERE is maximum [page1993average, fujita2018page]; therefore, $l$ is chosen to be $\lfloor N/2\rfloor$. To do that, we now compute the eigenvalues of the reduced density matrix on the subsystem u (or equivalently the subsystem d).
@@
where $\kappa>0$ is a constant, the ERE is lower bounded by $\Omega(N^{\kappa})$. Hence, by choosing $\kappa\to 0$, i.e., $\eta$ scales faster than $1/\sqrt{N}$, the corresponding lossy boson sampling is hard to simulate using our MPS method. Therefore, the approximability achieved by our MPS method exhibits a transition at $\eta=\Theta(1/\sqrt{N})$, which is consistent with the GBS case [oh2024classical].
@@
We now compare our results with those of previous studies [garcia2019simulating, oszmaniec2018classical]. In the previous work, classical algorithms are proposed whose classically simulable range of $\eta$ is $\eta=O(1/\sqrt{N})$. The algorithms employed a method that approximates the lossy input state as a state that is easy to classically simulate, such as the closest thermal state or the closest separable state. An obvious crucial limitation of this approach is that the closest thermal state or the closest separable state to a given lossy input state remains fixed once the parameters of the lossy boson sampling are determined. That means that when the system is not sufficiently lossy, no thermal state or separable state can approximate the lossy input state with adequate accuracy. Furthermore, the simulation accuracy cannot be improved by allocating additional computational resources because the closest easy-to-simulate state is fixed.
@@
Compared to that, although our MPS algorithm exhibits a similar scaling in an asymptotic regime as $\eta<O(1/\sqrt{N})$, i.e., our MPS method might not provide an advantage over the previously known classical algorithm asymptotically, it shows a wider simulable range than the previous work in a non-asymptotic regime, which is more practically relevant, because even if the loss rate is not high enough, our algorithm always runs and can achieve a target accuracy by using more computational resources, i.e., by increasing the bond dimension. We emphasize that this improvement significantly enhances simulation performance in simulating lossy GBS [qi2020regimes, oh2024classical, martinez2023classical]. To be more specific, Ref. [qi2020regimes] shows that the closest nonclassical state renders a similar classically simulable range of $\eta$, $\eta=O(1/\sqrt{N})$, which, however, was insufficient to simulate the experimental GBS. Nonetheless, Ref. [oh2024classical] developed a classical algorithm based on MPS, just like ours, and could simulate the state-of-the-art experimental GBS, which clearly indicates that the MPS method significantly outperforms other methods in practice even though the asymptotic scaling is similar.
@@
We now compare this with another previous result based on MPO, which numerically finds a similar classically simulable range $\eta=O(1/\sqrt{N})$ [oh2021classical]. Although the previous result yields the same strength in the sense that the accuracy can be improved by adding more computational resources, our work demonstrates a significant enhancement of the numerical simulability range. More specifically, the computational time cost of MPS and MPO simulations in a circuit of depth $D$ is $T=O(MDd^{4}\chi^{3})$ and $T=O(MDd^{8}\chi^{3})$, respectively. In terms of memory requirement, the memory cost of MPS is $O(\chi^{2}dM)$, while that of MPO is $O(\chi^{2}d^{2}M)$. In the MPO case, the local Hilbert space dimension $d$ is the square of that in the MPS case. This can lead to a higher computational cost when $d=N+1$ and $N$ is sufficiently large, as is the case in the boson sampling scenario. More crucially, as mentioned earlier, classical correlation can contribute to the increase in the bond dimension of MPO, whereas it cannot for MPS. Therefore, the bond dimensions between the two methods also differ between MPS and MPO. For example, the MPO method requires a bond dimension of the order of $10^{4}$, even for an input photon of less than $10$, with a loss rate of $0.7$. Our MPS method, on the other hand, requires only a bond dimension of the order of $10$ (note that the definitions of approximation errors differ, but a significant difference in bond dimension is evident).
@@
We also justify this commutation relation without the Haar-random assumption. It has been known that in the lossless case, the linear-optical circuit that renders the maximum entanglement entropy is the one that transforms the initial bosonic creation operators as [oh2024classical, oh2025classical]
@@ IV.4 Numerical analysis of the computational cost
To understand how large parameter regimes our classical algorithm can simulate in practice, we now directly estimate the bond dimension for simulating the boson sampling algorithm with the photon number $N$ and transmission rate $\eta$, with a fixed MPS approximation error $\epsilon$ of $0.01$. Our MPS approximation error is defined by the sum of the discarded eigenvalues when the system is bipartitioned into two halves. Before we present our results, we emphasize that for certain regimes, other existing methods, such as the Clifford-Clifford algorithm, may outperform our method [clifford2018classical, clifford2020faster]. Hence, one should interpret our results as a boundary set by our MPS method, rather than the boundary for the classically simulable regime for all possible methods.
In this numerical analysis, we consider three different families of circuits: (a) the worst-case construction, (b) the Haar-random unitary, and (c) an experimentally motivated structure. The purpose of this comparison is to examine how the required bond dimension varies across different choices of circuit unitaries. Here, the worst-case construction has $\theta_{j}=\pi/4$ for all $j$ to produce the largest amount of entanglement between the bipartitions we consider, and the experimentally motivated structure follows the same structure implemented in Ref. [zhong2020quantum]. For (c), although the experiment in the reference implements GBS, we constructed a circuit unitary from this setup. Since the number of modes of this unitary is restricted as $M=2k^{2}$ for an integer $k$, for a given photon number $N$, we set $k$ as the nearest integer to $N/\sqrt{2}$ so that $M=2k^{2}$ is the closest value to $N^{2}$. However, we inject $\lceil N/2\rceil$ photons into the middle of the modes of the first half, and the remaining $\lfloor N/2\rfloor$ photons into the middle of the modes of the second half. This is because the whole unitary in this case is constructed by combining two unitaries of half the size; inputting the same number of single photons in both bipartitions and inputting single photons in the middle can help maximize the entanglement.
We present the numerical results in Fig. 6. First of all, Fig. 6(a) presents the required bond dimension in the worst-case circuit. Fig. 6(b) is when the unitary of the boson sampling circuit is drawn from a Haar random. The required resource in this result is comparable to the worst-case scenario, indicating that highly entangled configurations are typical in Haar random cases. One can see that if the loss rate is approximately $50\%$, which is the case for GBS experiments in Refs. [zhong2021phase, deng2023gaussian, liu2025robust], and the input photon number is $40$, the required bond dimension is $10^{7}$, which is already quite large to implement in practice.
@@
Figure 6: Required bond dimension and memory to achieve an MPS approximation error $\epsilon=0.01$. Panels (a)–(c) show the required bond dimension, and panels (d)–(f) show the required memory. (a) Required bond dimension where the unitary is in the worst-case scenario. This implies the highest cost for a specific input photon number and photon loss. (b) Required bond dimension when the Haar random unitary is a unitary corresponding to the circuit in boson sampling. (c) Required bond dimension when a circuit unitary is the same structure as the GBS experiment of USTC [zhong2020quantum]. (d) Required memory where the unitary is in the worst-case scenario. (e) Required memory when the Haar random unitary is a unitary corresponding to the circuit in boson sampling. (f) Required memory when a circuit unitary is the same structure as the GBS experiment of USTC [zhong2020quantum].
According to the required bond dimension, the memory complexity of the circuit can be estimated as shown in Fig. 6(d)–(f) [oh2024classical]. Assuming that each complex number requires 8 bytes, the memory usage can be estimated as the total number of elements multiplied by 8 bytes:
@@ IV.6 Cat-state boson sampling
respectively. Note that as the amplitude of the cat state $\gamma\rightarrow 0$, the odd (even) cat state approaches $|1\rangle$ ($|0\rangle$), thereby recovering single-photon inputs. Moreover, cat state sampling is also known to be hard to classically simulate [rohde2013sampling]. The circuit has the same structure as before, except that the input of the first $N$ mode is an odd cat state. The photon loss can be gathered in front of the circuit as in Fig. 1, and the cat state input goes through the total photon loss channel and transforms to the following state,
@@ V.2 Entanglement Rényi entropy of the output state
Hence, it shows that our MPS method is efficient up to the point where the depth scales logarithmically with the number of qubits $n$. Notably, this noise threshold is consistent with a recent result in Ref. [nelson2024polynomial], which employs a completely different method.
Figure 9: Simulation of depolarizing noise in IQP circuit. (a) Pauli $Z$ noises can commute with the very first gate of the circuit, as it is diagonal in the computational basis. (b) The noises are sampled into Pauli gates, where each $X$ and $Y$ gate occurs with probability $q$. And the input $|+\rangle$ goes through the Pauli $Z$ noise channel and becomes a mixed state $\hat{\tau}$. (c) We sample from an ensemble decomposition of the mixed state, drawing $|\phi_{+}\rangle$ or $|\phi_{-}\rangle$ with equal probability $1/2$. The $Y$ gates can be decomposed into $X$ and $Z$ gates. Here, because IQP gates are diagonal in the $Z$ basis, we absorb all $Z$ gates into the ideal IQP layer; the shaded boundary indicates this absorption. In contrast, $X$ gates do not commute and therefore remain in the circuit.
@@ V.3 Generalization into depolarizing noise
Since the Rényi entropy of the state in a diagonal form depends only on the coefficients, the ERE of this state is exactly the same as that we obtained from the previous section. Therefore, as in the dephasing noise case, when the circuit depth $d\geq O(\log n/|\log(1-2q)|)=O(\log n/|\log(1-2p)|)$, the IQP sampling under depolarizing noise is efficiently classically simulable. When the noise rate is small, this depth approaches $O(\frac{1}{p}\log n)$, which corresponds to the critical depth in Ref. [nelson2024polynomial].
@@ V.4 Numerical analysis of noisy IQP sampling
Since we always assume the worst-case IQP circuit regardless of the circuit depth, even with $d=1$, a highly complicated IQP circuit is considered. Due to this reason, the required bond dimension is the maximum under the same number of qubits and the same noise rate per qubit and gate. Obviously, in practice, the complexity of an IQP circuit depends on its circuit depth [markov2008simulating, bremner2017achieving, oh2022classical, maslov2024fast]. Implementing a highly complicated IQP circuit in a very shallow depth is demanding in practice; thus, we expect that the bond dimension for extremely low-depth cases overestimates the complexity. In other words, for an extremely low depth, though the required bond dimension is shown to be high in our figure, the corresponding IQP sampling can be classically efficiently simulated because entanglement cannot spread to the overall qubits. Except for that case, when the circuit depth increases, the overall noise rate also increases; therefore, the complexity of the simulation decreases.
For IQP sampling, the previous work shows the simulability beyond the critical depth $d_{c}>\Theta(p^{-1}\log(p^{-1}))$ [rajakumar2025polynomial]. According to the reference, the phase transition for percolation occurs at depths $d_{c}\approx 33$ for noise rate $p=0.05$ and $d_{c}\approx 117$ for $p=0.02$. However, since the IQP circuit is classically simulable using MPS with a bond dimension of order $10^{4}$, which may still be possible using a large classical computer, our MPS algorithm can cover a much wider range; more specifically, ours can simulate an IQP circuit with lower depth than the critical depth of the previous work. For the calculation of simulable memory of IQP, we use the same memory resource formula as in boson sampling.
@@ VI Discussion and conclusion
We remark that although our method for lossy boson sampling covers Fock-state and cat-state inputs, whether it can be further generalized to an arbitrary input state is still open. Also, while we assumed a conventional IQP circuit [bremner2011classical], a recent study in Ref. [nelson2024polynomial] extended the IQP circuit to be augmented by CNOT gates. Our method does not immediately extend to those IQP circuits yet. It is open to improve our algorithm to cover a broader class of quantum circuits. Also, we assumed a lossy channel or Pauli noise for boson sampling and IQP sampling, respectively. Generalizing our method to cover more general noise channels would be an important open question.